OpenStudy (baileedoherty):

The distance from Phoenix to Tucson is 120 miles. A car made the journey in 1 hour and 45 minutes. What was the average speed of the car in miles per hour?

OpenStudy (photon336):

so the first thing you need to do is this: \[\frac{ distance }{ time } = speed \]

OpenStudy (photon336):

\[\frac{ distance }{ time } = \frac{ miles }{ hour }\]

OpenStudy (photon336):

so he traveled 120 miles in 1.75 hours. based on what I've shown you above how would you put this together?

OpenStudy (baileedoherty):

how did you get 1.75 for the time?

OpenStudy (frostyflame):

.75 is 3/4 of 1, so 3/4 of an hour is 45 minuets so what Photon is says is that you need to divide 120miles by the amount of hours which is 1.75

OpenStudy (baileedoherty):

I was able to get the correct answer after he explained it. now im just confused on the time part.

OpenStudy (frostyflame):

a hour is 60 minutes so 1/4 of an hour is 15 min and 3/4 is 45 minutes but 1/4 of an 1 is .25 so 3/4 is .75

OpenStudy (baileedoherty):

that makes more sense. but what would something like 2 hours and 3 minutes be?

OpenStudy (frostyflame):

so its still be 2 hours but the three minutes 1/20 of an hour

OpenStudy (baileedoherty):

how would I change it into a decimal?

OpenStudy (frostyflame):

it would be .05 and in a calculator type in 1/20 divided by 1

OpenStudy (baileedoherty):

okay thank you sorry its been awhile sense ive been in math

OpenStudy (frostyflame):

of course :)